OpenStudy (anonymous):

If x varies directly as y, and x = 35 when y=5, find x when y = 7.

OpenStudy (anonymous):

x varies directly as y so we can write x= k y

OpenStudy (anonymous):

x = 35 when y=5 so 35 = k*5 hence k= 7 now we have x= 7y so when y=7 x=7*7=49

OpenStudy (anonymous):

so x = 49 or ky?

OpenStudy (anonymous):

x = 49

OpenStudy (anonymous):

Awesome! Thank you!

OpenStudy (anonymous):

yw